OpenStudy (anonymous):

how to write a polynomial equation with given roots -1, 3±i, 2±3i

OpenStudy (dumbcow):

in general for these problems do these steps: -- assign each root equal to "x" -- work backwards to get one side to be zero -- the other side will be a "factor" of the polynomial

OpenStudy (dumbcow):

note: to get rid of the "i" you would square both sides

OpenStudy (anonymous):

let Q be that set of complex numbers, then $$\prod_{a\in Q}(x-a)$$ Is a polynomial with those roots.

OpenStudy (anonymous):

omg can you guys give me steps TAT i dont get anything from you guys right now

OpenStudy (dumbcow):

example \[x = 3\pm i\] \[x-3 = \pm i\] \[(x-3)^{2} = i^{2} =-1\] \[(x-3)^{2} +1 = 0\] \[x^{2}-6x+10 = 0\]

OpenStudy (anonymous):

oh god you are wow ok thanks so much i got this i got this

OpenStudy (dumbcow):

yw, make sure you multiply all the factors at the end to get polynomial with a "x^5" in it

OpenStudy (anonymous):

what did you say?

OpenStudy (dumbcow):

they gave you 5 roots, so the answer should be a 5th degree polynomial

OpenStudy (anonymous):

oh ok haha

OpenStudy (anonymous):

can you guys give me the answer so i can solve this multiple time

OpenStudy (dumbcow):

i thought you said you "got this" i will check your answer for you, if you do the rest of the work

OpenStudy (anonymous):

ok i will try

OpenStudy (dumbcow):

tell me what you get for the roots: 2+-3i

OpenStudy (anonymous):

x^2-10x+23

OpenStudy (dumbcow):

ok i have x^2 -4x +13 can you show your steps and ill tell you where you messed up

OpenStudy (anonymous):

oh sorry i type worng that was right

OpenStudy (anonymous):

i typed wrong one

OpenStudy (anonymous):

so whats next step

OpenStudy (anonymous):

i think i got the final answer can you check it

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

2x^3-12X^2+33x-23

OpenStudy (dumbcow):

no thats not even remotely close?? remember the 5 roots , 5th degree thing

OpenStudy (anonymous):

how can it form 5 roots...

OpenStudy (anonymous):

oh i got it

OpenStudy (anonymous):

wait teacher wait

OpenStudy (anonymous):

x^5-14x^4+57x^3-165x^2+248x-130

OpenStudy (anonymous):

i hope it does...

OpenStudy (dumbcow):

that looks better but coefficients are off, let me verify you started with this right \[(x+1)(x^{2}-6x+10)(x^{2}-4x+13)\]

OpenStudy (anonymous):

no... its x-1 what r u talking about....

OpenStudy (dumbcow):

-1 is the root \[x = -1\] \[x+1 = 0\] (x+1) is the factor

OpenStudy (anonymous):

so.. really. i gotta change that too? root thingy?

OpenStudy (dumbcow):

yeah

OpenStudy (anonymous):

ok i got the answer check it out

OpenStudy (anonymous):

x^5-8x^4+x^3-58X^2+12x+130

OpenStudy (dumbcow):

the "12x" is correct should be: \[x^{5}-9x^{4}+37x^{3}-71x^{2}+12x+130\]

OpenStudy (anonymous):

wow.............

OpenStudy (anonymous):

what did you multiply by last one

OpenStudy (dumbcow):

\[(x+1)(x^{2}-6x+10)(x^{2}-4x+13)\] \[(x^{3}-5x^{2}+4x+10)(x^{2}-4x+13)\]

OpenStudy (anonymous):

hey you solve wrong sigh.........................

OpenStudy (dumbcow):

??

OpenStudy (anonymous):

wtf you solved wrong..........

OpenStudy (dumbcow):

what do you mean i solved wrong? are you saying you disagree with the solution

OpenStudy (anonymous):

(x+1)(x2−6x+10) dont you multyply these

OpenStudy (dumbcow):

yes correct the next line shows the result

OpenStudy (anonymous):

its x^3-4x^2+4x+10.............. RU OK

OpenStudy (dumbcow):

when you expand you should have -6x^2 +x^2 = -5x^2

OpenStudy (anonymous):

how to expand,,, so tired sorry..

OpenStudy (dumbcow):

expand, multiply, distribute its all the same thing